Brent owns Beck Trucking. Seven years ago, he purchased a large-capacity dump truck for $115,000 to provide short-haul earth moving services. He sold it today for $45,000. Operating and maintenance costs averaged $9500 per year. A complete overhaul at the end of year 4 cost an extra $3200. (a) Calculate the annual cost of the truck at 7% per year. (b) If Brent estimates that Beck cleared at most $20,000 per year added revenue from using the truck, was the purchase economically advantageous?


1
Expert's answer
2021-11-03T10:37:48-0400

a

"AW=-115000(A/P,7\%,7)+45000(A/F,7\%,7)-9500-3200(A/F,7\%,4)\\=-115000(0.1855)+45000(0.1155)-9500-3200(0.2252)\\=-\$26,355.64"


b.

The annual worth of truck was -$26,355.64 while the revenue earned per year was less i.e., $20,000. So, the purchase was not economically advantageous.